Suppose eight workers died because of exposure to a toxic chemical. What is the YPLL for these workers, given their ages at death were 22, 25, 25, 30, 30, 47, 49, and 62 years? Use the age of 65 years as the endpoint. (Hint: YPLL = Predetermined end point age – Age of decedent who died prior to end point age).

  1. 230
  2. 26.8
  3. 214
  4. 74.3
Answer #1

Years of potential life lost can be calculated for the eight workers died because of exposure to a toxic chemical are as follows:

Here the predetermined end point age is given as 65 years old.

Thus, we can calculate each person's YPLL

  1. 1st person's age at death is 22 year old
  • YPLL of 1st persons is = Predetermined end point age - age of decedent who died prior to end point age

YPLL = 65 - 22 = 43

  • YPLL of 2nd persons is = Predetermined end point age - age of decedent who died prior to end point age

  YPLL = 65 - 25 = 40

  • YPLL of 3rd persons is = Predetermined end point age - age of decedent who died prior to end point age

YPLL = 65 - 25= 40

  • YPLL of 4th  persons is = Predetermined end point age - age of decedent who died prior to end point age

YPLL = 65 - 30 = 35

  • YPLL of 5th  persons is = Predetermined end point age - age of decedent who died prior to end point age

YPLL = 65 - 30 = 35

  • YPLL of 6th  persons is = Predetermined end point age - age of decedent who died prior to end point age

YPLL = 65 - 47 = 18

  • YPLL of 7th  persons is = Predetermined end point age - age of decedent who died prior to end point age

YPLL = 65 - 49 = 16

  • YPLL of 8th  persons is = Predetermined end point age - age of decedent who died prior to end point age

YPLL = 65 - 62 = 3

YPLL of a populatioin = sum of YPLL of each individuals

YPLL of 8 workers = 43+40+40+35+35+18+16+3 = 230

The answer is 230 (a)